# Non-Adaptive Multi-Stage Algorithm and Bounds for Group Testing with Prior Statistics

### v4

@@ V Numerical Evaluation
This section assesses the performance of the proposed MSGT algorithm by numerical study. First, in Subsection V-1, we provide a numerical evaluation to support our theoretical results and bounds given in Section III. Then, in Subsection V-2, we contrast the performance of MSGT with those of DD, ML, and MAP in a practical regime of $N$ and $K$. To generate the correlated prior information between adjacent items, we use the special GE process that reaches the infected state exactly $K$ times after $N$ steps, as described and analyzed in Section IV-D. The GE model is characterized by initial probabilities assigned to these two states, denoted as $\pi_{i}\in[0,1]^{2}$, as well as transition probabilities between them $\mathbf{\Phi}_{i}\in[0,1]^{2\times 2}$. These characteristics align well with the inputs required by the MSGT algorithm. In the practical scenarios tested (e.g., in the regime of COVID-19, when the test machine can simultaneously process a fixed small number of measurements [43, 44], or in sparse signal recovery in signal-processing with fixed vector size of input samples [9, 10, 11, 13, 64]), we show that the low computational complexity MSGT algorithm can reduce the number of pool tests by at least $25\%$.
@@
(b) Correlation matrix $\mathbf{\Phi}_{i}$ for $\tau=2$.
(c) Long memory priors - exact recovery.
Figure 7: (a) Number of computational operations in MSGT and MAP as given in Theorem 4 and Remark 1, respectively. (b) Example of correlation matrix $\mathbf{\Phi}_{i}$. (c) Probability of success of MSGT with exact prior statistics of 3-memory-steps Markov process, and with limited prior statistics assuming the Markov process has only one memory step. $N=1024,K=8$, 1000 iterations.
@@ V-1 Theoretical Analysis
In Fig. 4(a) we show the concentration of $\left|\mathcal{P}^{(S_{1,1})}\right|$, as obtained from the simulation, along with the bound on its expectation and on the worst-case that were calculated in Theorems 1 and 2 respectively. Note that the worst-case scenario regarding MSGT is when LVA filters the correct set of defective items. That may happen if the number of possibly defective items exceeds the threshold $\gamma K$. Since in MSGT, we only allow this deviation from the average and ignore the case of exceeding this threshold, our upper bound for the worst-case does not cover all potential realizations of $\left|\mathcal{P}^{(S_{1,1})}\right|$. Similarly, Fig. 4(b) demonstrates the concentration of $\left|\mathcal{P}^{(DD)}\right|$, as acquired through simulation, and the lower bound on its expectation as given in Theorem 3. Fig. 4(c) illustrates the numerically computed lower bound for $\gamma$, derived from the inequality provided in Proposition 1. For this simulation, we fix $N=500$, $K=3$, and calculate $\gamma$ value for a specific range of $T$ relative to the upper bound of ML. As explained in Section III, our conjecture asserts that any value of $\gamma$ surpassing this lower bound guarantees that MSGT performance will be at least on par with that of ML. Therefore, whenever computational resources allow, it is advisable to choose the value of $\gamma$ corresponding to the lower bound. This approach was followed in the subsequent simulations, and the practical outcomes presented in Subsection V-2 provide empirical support for our conjecture. Fig.7(a) compares the number of potential combinations to be examined in the MAP step, with and without the execution of the LVA step in MSGT. It can be observed that the LVA step performs an extensive filtering process, which allows MSGT to remain feasible even when executing MAP is no longer possible, especially in a regime below ML upper bound.
@@
The GE model considered in our numerical evaluations is a stationary Markov chain with $\tau=1$. Thus, $P\left(U_{i}\lvert U_{0},\ldots,U_{i-1}\right)=P\left(U_{i}\lvert U_{i-1}\right)$, for $i\in\{2,\ldots,N\}$. Substituting those priors, it follows that the converse of our problem is
@@
This bound is illustrated in the practical scenarios tested in Fig. 4 and Fig. 5. In Fig 8, we compare the lower and the upper bounds with GE model. The lower and upper bounds for GT with correlated prior statistics are given in (14) and in Section IV-D, respectively, while for GT without prior statistical information, the bounds are as given in [17].
Figure 8: Lower and upper bounds for GT with GE model (left). MSGT and corresponding MAP upper bound for GT with GE model (right). For MAP and ML upper bounds $\varepsilon=0.6$.
@@ V-2 Algorithm Evaluation
We demonstrate the performance of MSGT using simulation. The population is sampled from GE model, and the regime is $K=\mathcal{O}\left(N^{\theta}\right)$ with $\theta\leq 1/3$. The GE parameters serve as our prior statistics, but in practice, we ignore samples where the number of defective items does not match $K$. In addition, although in the complexity analysis, we considered Bernoulli encoder for simplification, here we use a near-constant column weight encoder that optimizes DND’s performance [46], with $\frac{\ln{2}}{K}T$ tests sampled randomly for each item. The parameter $\gamma$ was chosen to satisfy (3) and $L=500$ was chosen empirically. Fig. 5 compares MSGT to MAP, ML, and DD algorithms. We run DND and DD before ML and MAP for feasible runtime and memory consumption. The population includes $N\in\{500,1024,10000\}$ items and $K\in\{3,8,13\}$ defective items, respectively, and the empirical success probability is the average over $1000$ experiments. Note that for $N=1024,10000$ it is no longer possible to compare the performance since ML and MAP become infeasible (Fig. 5(b),5(c)).
Finally, MSGT was also tested with more complex probabilistic models. We sample the population based on a 3-memory-steps Markov process ($\mathbf{\Phi}_{i}$ is a $8\times 2$ matrix for all $i$). We execute MSGT using these prior statistics and also execute it with limited prior statistics, assuming that the process has only one memory step ($\mathbf{\Phi}_{i}$ is $2\times 2$). For example of $\mathbf{\Phi}_{i}$ representing a 2-memory-steps process, see Fig 7(b). The results are shown in Fig. 7(c). It is evident that utilizing the prior of long memory improves the success probability by $10\%$ in this scenario. We do note again that for the practical regime tested as in [55, 56, 8, 57], i.e., $N=1024$ and $K=8$, it is no longer possible to compare the performance since, unlike the efficient proposed MSGT algorithm, ML and MAP decoders for GT become infeasible.
@@ VI Conclusions and Future Work
In this work, we propose a MAP decoder for GT problems with arbitrary correlated sources. We analyze the performance of the MAP decoders for this case and provide, to the best of our knowledge, the only analysis that goes below $T_{ML}$, as the existing solutions in the literature do not offer an efficient and general solution for this regime. Our proposed MSGT can leverage the prior correlated information, within the LVA step, to significantly reduce the number of potential defective combinations. This approach yields an efficient computational solution that, as demonstrated in our simulation result for practical scenarios, can approach the minimum number of tests as MAP algorithm.
@@ Appendix A Recovery Algorithms
For the completeness of our proposed Algorithm 1, we present here all the algorithms that we use as integral components within MSGT, and provide detailed explanations for each.
@@ A-A Definitely Not Defective (DND)
DND algorithm [48, 49, 50, 19], which is shown in Algorithm 2 takes the testing matrix $\mathbf{X}$ and the results vector $Y$ as inputs. It systematically compares each column of the $\mathbf{X}$ with $Y^{T}$. If a specific column contains a value of $1$ while its corresponding test result is 00, we say that this column cannot explain the test result. Since each column represents one item in the population, if $\mathbf{X}(i)$ cannot explain $Y^{T}$, then DND marks this item as definitely non-defective. The output is the set of all such items, denoted by $\mathcal{P}^{(DND)}$.
@@ A-B Definite Defectives (DD)
Algorithm 3 Definite Defective (DD) [19]
@@ A-C Priors Update
The algorithm translates the information gathered during the first stage of MSGT into the state space. In particular, if an element is detected as defective, its probability of being defective is updated to $1$. Additionally, the transition probabilities are modified so that the transition to the defective state becomes deterministic.
@@
Input: $\{\mathbf{\pi}_{i},\mathbf{\Phi}_{i}\}_{i=1}^{N},\mathcal{P}^{(S_{1,2})},\mathcal{P}^{(DD)}$
Output: $\{\mathbf{\pi}_{i},\mathbf{\Phi}_{i}\}_{i=1}^{N}$
1:for each $i\in\mathcal{P}^{(\mathrm{DD})}$ do
2:    $\pi_{i}\leftarrow 0$
3:    if $i\neq 1$ then
4:         $\boldsymbol{\Phi}_{i-1}[:,0]\leftarrow\boldsymbol{0}$
5:         $\boldsymbol{\Phi}_{i-1}[:,1]\leftarrow\boldsymbol{1}$
6:    end if
7:end for
8:return $\{\mathbf{\pi}_{i},\mathbf{\Phi}_{i}\}_{i=1}^{N}$
@@ A-D List Viterbi Algorithm (LVA)
Our variations of the LVA algorithm [32], outlined in Algorithm 5, is designed to return the $L$ most likely sequences, representing the estimation of the whole population, for a given $L\geq 1$. In the algorithm we suggest herein, the key differences are: (1) the population sequence in GT replaces the time sequence in classical LVA as given in [32], and (2) we use the aggregated sequence likelihood instead of the general cost function presented in the original paper. In particular, in the suggested algorithm, as we traverse the trellis diagram, we are iteratively maximizing the likelihood of the sequence representing the status of the population, tested with correlated and non-uniform prior statistics, denoted by $\Psi$ (see line 15).
@@
Input: $L,\tau,\left\{\mathbf{\pi}_{i}\right\}_{i=1}^{N},\{\mathbf{\Phi}_{i}\}_{i=1}^{N}$
@@
3:    $\Psi(1,s,1)\leftarrow\left\{\mathbf{\pi}_{i}\right\}_{i=1}^{N}\left(s\right)$triangleright initial probability for each
                                     state
4:    $\xi\left(n,s,l\right)\leftarrow s$triangleright previous state of each state
5:    for $l\leftarrow 2$ to $L$ do
6:         $\Psi(1,s,l)\leftarrow 0$
7:         $\xi\left(n,s,l\right)\leftarrow s$
8:    end for
9:end for
10:Recursion:
11:for $n\leftarrow 2$ to $N$ do
12:    for $s_{2}\leftarrow 1$ to $2^{\tau}$ do
13:         for $l\leftarrow 1$ to $L$ do
14:             $\triangleright$ find the probability of the $l$-most likely
                   previous state and its rank
15:             $\Psi(n,s_{2},l)=\max^{(l)}_{{\begin{subarray}{c}s_{1}\in\{1,...,2^{\tau}\}\\ k\in\{1,..,L\}\end{subarray}}}\newline \hskip 103.50017pt\left\{\Psi(n-1,s_{1},k)\Phi_{n}\left(s_{1},s_{2}\right)\right\}$
16:             $\left({s_{1}}^{*},k^{*}\right)=\arg\max^{(l)}_{{\begin{subarray}{c}s_{1}\in\{1,...,2^{\tau}\}\\ k\in\{1,..,L\}\end{subarray}}}\newline \hskip 103.50017pt\left\{\Psi(n-1,s_{1},k)\Phi_{n}\left(s_{1},s_{2}\right)\right\}$
17:             $\xi\left(n,s_{2},l\right)\leftarrow{s_{1}}^{*}$
18:             $\chi\left(n,l\right)\leftarrow k^{*}$
@@
24:    $\mathbf{Z}_{l,N}\leftarrow\arg\max_{s\in\{1,...,2^{\tau}\}}\left\{\Psi(N,s,l)\right\}$
25:    $l_{N}\leftarrow\chi\left(n,l\right)$
26:    for $n\leftarrow N-1$ to $1$ do
27:         $\mathbf{Z}_{l,n}\leftarrow\xi\left(n+1,\mathbf{Z}_{l_{n+1},n+1},l_{n+1}\right)$
28:         $l_{n}\leftarrow\chi\left(n+1,l_{n+1}\right)$
@@
In the recursion stage, we iterate from the second item to the last, and for each item, compute all the $2^{\tau}\times 2^{\tau}$ possible transition probabilities between all the possible states he could be in and the states of its predecessor. For each item, state, and rank, we set in $\Psi$ the $L$-most likely probabilities of the overall sequences from the first item until the current item and state (line 15), and we set in $\xi$ the corresponding previous state and in $\chi$ the corresponding rank (lines 17-18). Here, $\max^{(l)}$ denotes the $l$-largest value.
@@
A-E Maximum A Posteriori (MAP)
The MAP estimator [53], as shown in Algorithm 6, returns the set $c\in\mathbf{C}$ with the highest maximum a posteriori probability among all the sets, that is, the set $c$ that obtains the maximum $P\left(Y\lvert\mathbf{X},c\right)P\left(c\right)$. In this expression, the first probability represents the likelihood of obtaining the results $Y$ from a group test using the testing matrix $\mathbf{X}$ and the given set $c$ as the defective set. If the set $c$ and the given testing matrix $\mathbf{X}$ cannot explain for $Y$, then this probability is equal to zero. The second probability corresponds to the prior probability of the overall defective set $c$, calculated using $\left\{\mathbf{\pi}_{i}\right\}_{i=1}^{N},\left\{\mathbf{\Phi}_{i}\right\}_{i=1}^{N}$. The MAP algorithm returns the estimated defective set, denoted by $\hat{\mathcal{K}}$.

### v5

@@ V Numerical Evaluation
This section assesses the performance of the proposed 2SDGT algorithm by numerical study. First, in Subsection V-1, we provide a numerical evaluation to support our theoretical results and bounds given in Section III. Then, in Subsection V-2, we contrast the performance of 2SDGT with those of DD, ML, and MAP in a practical regime of $N$ and $K$. Although Section II adopts a random-$K$ model (with a design interval $(K_{\min},K_{\max})$), the simulations in this section are conducted for a fixed realized value of $K$ for comparison with the existing solutions in the GT literature [8, 4, 3]. This is also consistent with the theory, since the analytical results can be applied conditionally on the event $\{K=k\}$ for a typical value $k\in[K_{\min},K_{\max}]$, while the probability of atypical realizations $\{K\notin[K_{\min},K_{\max}]\}$ is controlled separately via exponential tail bounds (Chernoff-type concentration) as discussed in Section II and Section III-C. To generate the correlated prior information between adjacent items, we use the special Gilbert-Elliott process that reaches the infected state exactly $K$ times after $N$ steps, as described and analyzed in Section IV-D. The Gilbert-Elliott model is characterized by initial probabilities assigned to these two states, denoted as $\pi_{i}\in[0,1]^{2}$, as well as transition probabilities between them $\mathbf{\Phi}_{i}\in[0,1]^{2\times 2}$. These characteristics align well with the inputs required by the 2SDGT algorithm. In the practical scenarios tested (e.g., in the regime of COVID-19, when the test machine can simultaneously process a fixed small number of measurements [44, 9], or in sparse signal recovery in signal-processing with fixed vector size of input samples [32, 63, 1, 26, 42]), we show that the low computational complexity 2SDGT algorithm can reduce the number of pooled tests by at least $25\%$.
@@
(b) Long memory priors - exact recovery.
Figure 8: (a) Number of computational operations in 2SDGT and MAP as given in Theorem 4 and Remark 4, respectively. (b) Probability of success of 2SDGT with exact prior statistics of 3-memory-steps Markov process, and with limited prior statistics assuming the Markov process has only one memory step. $N=1024,K=8$, 1000 iterations.
@@ V-1 Theoretical Analysis
In Fig. 5(a) we show the concentration of $\left|\mathcal{P}^{(S_{1,1})}\right|$, as obtained from the simulation, along with the bound on its expectation and the upper deviation bound that were calculated in Theorems 1 and 2 respectively. Under the random-$K$ model, these theoretical bounds are interpreted conditionally on the realized value of $K$ (as done here), while the probability that $K$ falls outside the design interval $(K_{\min},K_{\max})$ is negligible by concentration and is accounted for separately in the overall error analysis. Note that the worst-case scenario regarding 2SDGT is when LVA filters the correct set of defective items. That may happen if the number of possibly defective items exceeds the threshold $\gamma K$. Since in 2SDGT, we only allow this deviation from the average and ignore the case of exceeding this threshold, our upper bound for the worst-case does not cover all potential realizations of $\left|\mathcal{P}^{(S_{1,1})}\right|$. Similarly, Fig. 5(b) demonstrates the concentration of $\left|\mathcal{P}^{(DD)}\right|$, as acquired through simulation, and the lower bound on its expectation as given in Theorem 3. Fig. 5(c) illustrates the numerically computed lower bound for $\gamma$, derived from the inequality provided in Proposition 1. For this simulation, we fix $N=500$, $K=3$, and calculate $\gamma$ value for a specific range of $T$ relative to the upper bound of ML. Fig.8(a) compares the number of potential combinations to be examined in the MAP step, with and without the execution of the LVA step in 2SDGT. It can be observed that the LVA step performs an extensive filtering process, which allows 2SDGT to remain feasible even when executing MAP is no longer possible, especially in a regime below ML upper bound.
@@
The Gilbert-Elliott model considered in our numerical evaluations is a stationary Markov chain with $\tau=1$. Thus,
$P\left(U_{i}\lvert U_{0},\ldots,U_{i-1}\right)=P\left(U_{i}\lvert U_{i-1}\right),\quad\text{for}\quad i\in\{2,\ldots,N\}.$
Substituting those priors, it follows that the converse of our problem is
@@
This bound is illustrated in the practical scenarios tested in Fig. 5 and Fig. 6. In Fig 9, we compare the lower and the upper bounds with Gilbert-Elliott model. The lower and upper bounds for GT with correlated prior statistics are given in (17) and in Section IV-D, respectively, while for GT without prior statistical information, the bounds are as given in [8].
Figure 9: Lower and upper bounds for GT with GE model (left). 2SDGT and corresponding MAP upper bound for GT with GE model (right). For MAP and ML upper bounds $\varepsilon=0.6$.
@@ V-2 Algorithm Evaluation
We demonstrate the performance of 2SDGT using simulation. The population is sampled from Gilbert-Elliott model, and the regime is $\mu=\mathcal{O}\left(N^{\theta}\right)$ with $\theta\leq 1/3$. The Gilbert-Elliott parameters serve as our prior statistics, but in practice, we ignore samples where the number of defective items does not match $K$. Equivalently, the reported success probabilities are conditional on a fixed realized $K$, which matches the way the analytical guarantees are applied on the typical event $K\in[K_{\min},K_{\max}]$ in the random-$K$ formulation. In addition, although in the complexity analysis, we considered Bernoulli encoder for simplification, here we use a near-constant column weight encoder that optimizes DND’s performance [37], with $\frac{\ln{2}}{K}T$ tests sampled randomly for each item. In all experiments, the parameter $\gamma$ is selected to satisfy a target upper bound $\eta$ on the probability that the number of surviving candidates after Stage 1 exceeds $\gamma K_{\max}$ (see (9)), and $L=500$ was chosen empirically. We fix $\eta=10^{-2}$, which was found to provide a good tradeoff between computational complexity and recovery performance across all tested regimes. The resulting value of $\gamma$ is computed according to the analytical bounds derived in Section III-B and is held fixed across Monte Carlo trials. Fig. 6 compares 2SDGT to MAP, ML, DD, and SCOMP (an algorithm that starts with DD and sequentially marks extra items as defective, ensuring the result is a satisfying set [3]) algorithms. We run DND and DD before ML and MAP for feasible runtime and memory consumption. The population includes $N\in\{500,1024,10000\}$ items and $K\in\{3,8,13\}$ defective items, respectively, and the empirical success probability is the average over $1000$ experiments. Note that for $N=1024,10000$ it is no longer possible to compare the performance since ML and MAP become infeasible (Fig. 6(b),6(c)).
Finally, 2SDGT was also tested with more complex probabilistic models. We sample the population based on a 3-memory-steps Markov process ($\mathbf{\Phi}_{i}$ is a $8\times 2$ matrix for all $i$). We execute 2SDGT using these prior statistics and also execute it with limited prior statistics, assuming that the process has only one memory step ($\mathbf{\Phi}_{i}$ is $2\times 2$). For example of $\mathbf{\Phi}_{i}$ representing a 2-memory-steps process, see Fig 1. The results are shown in Fig. 8(b). It is evident that utilizing the prior of long memory improves the success probability by $10\%$ in this scenario. We do note again that for the practical regime tested as in [57, 45, 20, 36], i.e., $N=1024$ and $K=8$, it is no longer possible to compare the performance since, unlike the efficient proposed 2SDGT algorithm, ML and MAP decoders for GT become infeasible.
@@ VI Conclusions and Future Work
In this work, we propose a MAP decoder for GT problems with general correlated priors that admit a trellis representation. We analyze the performance of the MAP decoders for this case and provide, to the best of our knowledge, the only analysis that goes below $T_{ML}$, as the existing solutions in the literature do not offer an efficient and general solution for this regime. Our proposed 2SDGT can leverage the prior correlated information, within the LVA step, to significantly reduce the number of potential defective combinations. This approach yields an efficient computational solution that, as demonstrated in our simulation result for practical scenarios, can approach the minimum number of tests as MAP algorithm.
@@ Appendix A Recovery Algorithms
For the completeness of our proposed Algorithm 1, we present here all the algorithms that we use as integral components within 2SDGT, and provide detailed explanations for each.
@@ A-A Definitely Not Defective (DND)
DND algorithm [38, 16, 15, 3], which is shown in Algorithm 2 takes the testing matrix $\mathbf{X}$ and the results vector $Y$ as inputs. It systematically compares each column of the $\mathbf{X}$ with $Y^{T}$. If a specific column contains a value of $1$ while its corresponding test result is 0, we say that this column cannot explain the test result. Since each column represents one item in the population, if $\mathbf{X}(i)$ cannot explain $Y^{T}$, then DND marks this item as definitely non-defective. The output is the set of all such items, denoted by $\mathcal{P}^{(DND)}$.
@@ A-B Definite Defectives (DD)
Algorithm 3 Definite Defective ($D$) [3]
@@ A-C Priors Update
The updatePriors procedure incorporates the information obtained in Stage 1 into the prior model. When Stage 1 determines that an item is certainly defective or non-defective, we treat this as fixed information and update the prior accordingly. If the item lies in the first $\tau$ positions, we modify the initial distribution so that it assigns probability only to configurations consistent with the known value. If the item appears at a later position, we restrict the transition probabilities to this item to the detected state only.
@@
Input: $\mathbf{\pi},\{\mathbf{\Phi}_{i}\}_{i=\tau+1}^{N},\mathcal{P}^{(S_{1,2})},\mathcal{P}^{(DD)}$
Output: $\mathbf{\pi},\{\mathbf{\Phi}_{i}\}_{i=\tau+1}^{N}$
1:for each $i\in\mathcal{P}^{(S_{1,2})}\cup\mathcal{P}^{(DD)}$ do
2:  Let $u^{\star}\leftarrow 0$ if $i\in\mathcal{P}^{(S_{1,2})}$, otherwise $u^{\star}\leftarrow 1$.
3:  if $i\leq\tau$ then
$\pi^{\prime}(u_{1},\ldots,u_{\tau})\leftarrow\frac{\pi(u_{1},\ldots,u_{\tau})\,\mathbf{1}\{u_{i}=u^{\star}\}}{\sum_{(v_{1},\ldots,v_{\tau})}\pi(v_{1},\ldots,v_{\tau})\,\mathbf{1}\{v_{i}=u^{\star}\}}.$
4:  else
$\Phi^{\prime}_{i}(u_{i}\mid u_{i-1},\ldots,u_{i-\tau})\leftarrow\begin{cases}1,&u_{i}=u^{\star},\\ 0,&u_{i}\neq u^{\star},\end{cases}$
5:  end if
6:end for
@@ A-D List Viterbi Algorithm (LVA)
Our variations of the LVA algorithm [56], outlined in Algorithm 5, is designed to return the $L$ most likely sequences, representing the estimation of the whole population, for a given $L\geq 1$. In the algorithm we suggest herein, the key differences are: (1) the population sequence in GT replaces the time sequence in classical LVA as given in [56], and (2) we use the aggregated sequence likelihood instead of the general cost function presented in the original paper. In particular, in the suggested algorithm, as we traverse the trellis diagram, we are iteratively maximizing the likelihood of the sequence representing the status of the population, tested with correlated and non-uniform prior statistics, denoted by $\Psi$ (see line 16).
@@
Input: $L,\tau,{\color[rgb]{0,0,0}\mathbf{\pi},\{\mathbf{\Phi}_{i}\}_{i=\tau+1}^{N}}$
@@
3:  $\color[rgb]{0,0,0}\Psi(\tau,s,1)\leftarrow\pi(s)$} $\triangleright$ initial probability for each state
4:  ${\color[rgb]{0,0,0}\xi(\tau,s,1)}\leftarrow s$ $\triangleright$ previous state of each state
5:  $\color[rgb]{0,0,0}\chi(\tau,s,1)\leftarrow 1$}
6:  for $l\leftarrow 2$ to $L$ do
7:    $\Psi(\tau,s,l)\leftarrow 0$
8:    ${\color[rgb]{0,0,0}\xi(\tau,s,l)}\leftarrow s$
9:    $\chi(\tau,s,l)\leftarrow 1$
10:  end for
11:end for
12:Recursion:
13:for $n\leftarrow\tau+1$ to NN do
14:  for $s_{2}\leftarrow 1$ to $2^{\tau}$ do
15:    for $l\leftarrow 1$ to $L$ do
16:
$\Psi(n,s_{2},l)\;=\;\max\nolimits^{(l)}_{\begin{subarray}{c}s_{1}\in\{1,\dots,2^{\tau}\}\\ k\in\{1,\dots,L\}\end{subarray}}\Big\{\Psi(n-1,s_{1},k)\;\Phi_{n}(s_{1},s_{2})\Big\}$
$(\,s_{1}^{\ast},\;k^{\ast}\,)\;=\;\arg\max\nolimits^{(l)}_{\begin{subarray}{c}s_{1}\in\{1,\dots,2^{\tau}\}\\ k\in\{1,\dots,L\}\end{subarray}}\Big\{\Psi(n-1,s_{1},k)\;\Phi_{n}(s_{1},s_{2})\Big\}$
17:     $\xi(n,s_{2},l)\leftarrow s_{1}^{\ast}$
18:     $\chi(n,s_{2},l)\leftarrow k^{\ast}$
@@
24:  $Z_{l,N}\leftarrow\arg\max_{s}\Psi(N,s,l)$
25:  $l_{N}\leftarrow{\color[rgb]{0,0,0}l}$
26:  for $n\leftarrow N-1$ to $\tau$ do
27:    $Z_{l,n}\leftarrow\xi\big(n+1,\;Z_{l,n+1},\;l_{n+1}\big)$
28:    $l_{n}\leftarrow\chi\big(n+1,\;Z_{l,n+1},\;l_{n+1}\big)$
@@
In the recursion stage, we iterate from the second item to the last, and for each item, compute all the $2^{\tau}\times 2^{\tau}$ possible transition probabilities between all the possible states he could be in and the states of its predecessor. For each item, state, and rank, we set in $\Psi$ the $L$-most likely probabilities of the overall sequences from the first item until the current item and state (line 16), and we set in $\xi$ the corresponding previous state and in $\chi$ the corresponding rank (lines 17-18). Here, $\max^{(l)}$ denotes the $l$-largest value.
@@
A-E Candidate Enumeration via getAllCombinations
In Stage 2 of 2SDGT (Algorithm 1), after obtaining a list of candidate population trajectories from the LVA, we must extract candidate defective sets to be evaluated by the MAP estimator. This task is performed by the procedure getAllCombinations, as shown in Algorithm 6. Given a set of indices $\mathbf{V}^{(l)}\subseteq\mathcal{N}$ corresponding to items estimated as defective in the $l$-th LVA trajectory, the goal of getAllCombinations is to enumerate all subsets whose cardinality lies within a prescribed range. This enables 2SDGT to handle uncertainty in the true number of defectives while keeping the candidate list size computationally feasible.
Under the random-$K$ model in Section II, the algorithm does not assume knowledge of the realized $K$. Instead, Stage 2 operates with a design interval $(K_{\min},K_{\max})$, and getAllCombinations enumerates all subsets whose cardinality satisfies $K_{\min}\;\leq\;|c|\;\leq\;K_{\max},c\subseteq\mathbf{V}^{(l)}$. This modification ensures robustness to fluctuations in the realized number of defectives while preserving correctness on the typical event $\{K\in[K_{\min},K_{\max}]\}$. We note that the procedure is purely combinatorial and does not rely on probabilistic assumptions and that the upper bound ${\color[rgb]{0,0,0}\lfloor}\gamma K_{\max}{\color[rgb]{0,0,0}\rfloor}$ ensures that the total number of candidate sets remains manageable, as analyzed in Section III. In practice, if $|\mathbf{V}^{(l)}|<K_{\min}$, the procedure returns the empty set, and the corresponding trajectory is discarded.
Algorithm 6 getAllCombinations
Input: $\mathbf{V}$, $K_{\min}$, $K_{\max}$
Output: $\mathcal{C}$
[…]
2:for $k=K_{\min}$ to $K_{\max}$ do
3:  for each subset $c\subseteq\mathbf{V}$ such that $|c|=k$ do
4:    $\mathcal{C}\leftarrow\mathcal{C}\cup\{c\}$
5:  end for
6:end for
7:return $\mathcal{C}$
A-F Maximum A Posteriori (MAP)
The MAP estimator [29], as shown in Algorithm 7, returns the set $c\in\mathbf{C}$ with the highest maximum a posteriori probability among all the sets, that is, the set $c$ that obtains the maximum $P\left(Y\lvert\mathbf{X},c\right)P\left(c\right)$. In this expression, the first probability represents the likelihood of obtaining the results $Y$ from a group test using the testing matrix $\mathbf{X}$ and the given set $c$ as the defective set. If the set $c$ and the given testing matrix $\mathbf{X}$ cannot explain for $Y$, then this probability is equal to zero. The second probability corresponds to the prior probability of the overall defective set $c$, calculated using $\color[rgb]{0,0,0}\mathbf{\pi},\{\mathbf{\Phi}_{i}\}_{i=\tau+1}^{N}$}. The MAP algorithm returns the estimated defective set, denoted by $\hat{\mathcal{K}}$.